# High-probability Convergence Guarantees of Decentralized SGD

### v3

@@ Appendix E Proofs for Non-convex Costs
    Unknown time horizon. In this case, we choose a time-varying step-size $\alpha_{t}=\frac{C^{\prime}}{\sqrt{t+1}}$, for all $t\geq 1$ and $C^{\prime}=\sqrt{2}C$, where $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{6\sigma\sqrt{10dL}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$, again guaranteeing that all the step-size conditions are satisfied. Noting that $\alpha_{t}\leq C$, we get $\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}\leq\frac{C}{(1-\lambda^{2})}$, hence, with probability at least $1-\delta$
@@
    $+\frac{3840\sqrt{2}\sigma^{2}dLC\log(T+1)}{n\sqrt{T+1}}+\frac{384\lambda^{2}L^{2}(\sigma^{2}+A^{2})C^{2}}{(1-\lambda)^{2}\sqrt{T+1}}.$
    Note that in the unknown time horizon case we lose the linear speed-up in the number of users, as, even in the case $C=\frac{\sqrt{n}}{6\sigma\sqrt{10dL}}$, we have
    $\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{dL}\big(\Delta_{f}+\log(\nicefrac{{T+1}}{{\delta}})\big)}{\sqrt{n(T+1)}}$
    $+\frac{\sigma\sqrt{d}\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}d(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
@@
Appendix F Proofs for Strongly Convex Costs
@@
where in $(i)$ we used the fact that $(1+\frac{a}{2})(1-a)\leq(1-\frac{a}{2})$ for any $a>0$, while $(ii)$ follows by setting $\nu\leq\frac{\mu}{24a\overline{\sigma}^{2}K}$ and from the step-size choice $\alpha_{t}\leq\frac{1}{\mu}$. Plugging (F) in (F), using the shorthand $E_{t}=\alpha_{t}^{2}n\sigma^{2}+9\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu$ and taking the full expectation, we get
@@
where $(i)$ follows from Proposition 3, in $(ii)$ we set $\theta=\frac{\widetilde{\lambda}}{2}$, while $(iii)$ follows from the fact that $\|\nabla\mathbf{f}^{t}\|^{2}\leq 2L^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|+2\|\nabla\mathbf{f}^{\star}\|^{2}$ (recall Proposition 7). Choosing $\alpha_{t}\leq\frac{\sqrt{1-\lambda}}{2\overline{\sigma}\lambda\sqrt{6(t+t_{0}+2)K}}$ and plugging (F) in (F), we get
@@
where we used Proposition 1 in the first and Lemma 11 in the second inequality, with $D_{t}=\frac{5\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{t}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(t+t_{0})^{a\mu}}\Big)$. Similarly, we use (F), to get
@@
Theorem 3.7.
Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{17280d\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
@@
$\times\exp\bigg(\frac{3\alpha_{t}^{2}\nu^{2}n(t+t_{0}+1)^{2}\sigma^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{2}+120\alpha_{t}^{2}\nu\sigma^{2}d(t+t_{0}+1)L\bigg)$
@@
where in $(a)$ we used the fact that $F_{t}$ and $\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}$ are $\mathcal{F}_{t}$-measurable, $(b)$ follows from Proposition 4, in $(c)$ we use Lemma 2, Proposition 1 and impose the condition $\alpha_{t}\leq\frac{1}{4\sigma\sqrt{15(t+t_{0}+1)dL}}$, while $(d)$ follows from Proposition 7 and the definition of $F_{t}$, with $b_{t}=\Big(1-\alpha_{t}\mu+3\alpha_{t}^{2}\nu(t+t_{0}+1)L\Big)\frac{t+t_{0}+1}{t+t_{0}}$, $c_{t}=\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}$ and $d_{t}=120\alpha_{t}^{2}\sigma^{2}d(t+t_{0}+1)L$. Taking the full expectation and applying Proposition 4, we get
@@
where the first inequality follows from $\alpha_{t}\leq\frac{1}{\mu}$. Using (40) and (41) in (F), it follows that
@@
where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$. Define $G_{1}\coloneqq\frac{8640\kappa\sigma^{2}d}{\mu}$, $G_{2}\coloneqq\frac{1296\kappa^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)}$, $G_{4}\coloneqq\frac{6480\kappa^{3}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and $G_{3}\coloneqq\frac{373248n\kappa^{4}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}}$, and plug into (F), to get
@@
Using (F), it can be readily verified that choosing
$\epsilon_{t}^{1}$ $=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})+4G_{1}}{n(t+t_{0})}+\frac{4G_{2}}{n(t+t_{0})^{2}}+\frac{(t_{0}+2)^{3}\Delta_{f}+4G_{3}\log(t+t_{0})+4G_{4}/5(t_{0}+1)^{5}}{n(t+t_{0})^{3}},$ (44)
@@
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg($ $\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}d\kappa/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}$
$+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
@@
Appendix G Numerical Experiments
In this section we provide some numerical results, which serve to verify our theory.
@@ Appendix G Numerical Experiments
We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive semi-definite, making each $f_{i}$ strongly convex, with unbounded gradients. We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [73]. When queried by user $i\in[n]$ in iteration $t\geq 1$, the $\mathcal{SFO}$ returns a noisy gradient of $f_{i}$ evaluated at $x^{t}_{i}$, i.e., $g^{t}_{i}=A_{i}x^{t}_{i}+b_{i}+z^{t}_{i}$, where $z^{t}_{i}\in\mathbb{R}^{d}$ is a zero-mean Gaussian random vector, making the noise consistent with assumption (A4). We use the DSGD method outlined in Algorithm 1, with the time-varying step-size schedule $\alpha_{t}=\frac{1}{t+1}$ and shared initializatio $x^{1}_{i}=0$, for all $i\in[n]$. We are interested in testing the following two facets of our theory.
@@
Appendix H Deriving the Transient Time

### v5

@@ Appendix F Proofs for Non-convex Costs
    Unknown time horizon. In this case, we choose a time-varying step-size $\alpha_{t}=\frac{C^{\prime}}{\sqrt{t+1}}$, for all $t\geq 1$ and $C^{\prime}=\sqrt{2}C$, where $C=\min\left\{\frac{1}{2L},\frac{n}{9\sigma^{2}},\frac{1-\lambda}{\lambda LB\sqrt{48}},\frac{\sqrt{n}}{3\sigma\sqrt{5L}},\frac{\sqrt[3]{n}(1-\lambda)^{2/3}}{\overline{\sigma}^{2/3}\lambda^{2}L^{2/3}\sqrt[3]{9}}\right\}$, again guaranteeing that all the step-size conditions are satisfied. Noting that $\alpha_{t}\leq C$, we get $\sum_{t\in[T]}\alpha_{t}\lambda^{2(t-1)}\leq\frac{C}{(1-\lambda^{2})}$, hence, with probability at least $1-\delta$
@@
    $+\frac{240\sqrt{2}\sigma^{2}LC\log(T+1)\big(1+\log(2d)\big)}{n\sqrt{T+1}}+\frac{384\lambda^{2}L^{2}(\sigma^{2}+A^{2})C^{2}}{(1-\lambda)^{2}\sqrt{T+1}}.$
    Note that in the unknown time horizon case we lose the linear speed-up in the number of users, as, even in the case $C=\frac{\sqrt{n}}{3\sigma\sqrt{5L}}$, we have
    $\frac{1}{nT}$ $\sum_{t\in[T]}\sum_{i\in[n]}\|\nabla f(x^{t}_{i})\|^{2}=\mathcal{O}\Bigg(\frac{\sigma\sqrt{L}\big(\Delta_{f}+\log(\nicefrac{{2dT}}{{\delta}})\big)}{\sqrt{n(T+1)}}$
    $+\frac{\sigma\Delta_{x}L^{3/2}}{(1-\lambda^{2})\sqrt{n(T+1)}}+\frac{n\lambda^{2}L(\sigma^{2}+A^{2})}{\sigma^{2}(1-\lambda)^{2}\sqrt{T+1}}\Bigg),$
@@
Appendix G Proofs for Strongly Convex Costs
@@
where in $(i)$ we used the fact that $(1+\frac{a}{2})(1-a)\leq(1-\frac{a}{2})$ for any $a>0$, while $(ii)$ follows by setting $\nu\leq\frac{\mu}{24a\overline{\sigma}^{2}K}$ and from the step-size choice $\alpha_{t}\leq\frac{1}{\mu}$. Plugging (G) in (G), using the shorthand $E_{t}=\alpha_{t}^{2}n\sigma^{2}+9\alpha_{t}\|\nabla\mathbf{f}^{\star}\|^{2}/2\mu$ and taking the full expectation, we get
@@
where $(i)$ follows from Proposition 3, in $(ii)$ we set $\theta=\frac{\widetilde{\lambda}}{2}$, while $(iii)$ follows from the fact that $\|\nabla\mathbf{f}^{t}\|^{2}\leq 2L^{2}\|{\mathbf{x}^{t}}-{\mathbf{x}^{\star}}\|+2\|\nabla\mathbf{f}^{\star}\|^{2}$ (recall Proposition 7). Choosing $\alpha_{t}\leq\frac{\sqrt{1-\lambda}}{2\overline{\sigma}\lambda\sqrt{6(t+t_{0}+2)K}}$ and plugging (G) in (G), we get
@@
where we used Proposition 1 in the first and Lemma 11 in the second inequality, with $D_{t}=\frac{5\alpha_{t}^{2}\lambda^{2}L^{2}}{1-\lambda}\Big(\frac{4an\sigma^{2}\alpha_{t}}{a\mu-1}+\frac{9\|\nabla\mathbf{f}^{\star}\|^{2}}{\mu^{2}}+\frac{(1+t_{0})^{a\mu}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(t+t_{0})^{a\mu}}\Big)$. Similarly, we use (G), to get
@@
Theorem 2.
Let (A1)-(A4) and (A6) hold, let the step-size be given by $\alpha_{t}=\frac{a}{t+t_{0}}$ and let $x^{1}_{i}=x^{1}_{j}$, for all $i,j\in[n]$. If $a=\frac{6}{\mu}$ and $t_{0}\geq\max\left\{6,\frac{1960\sigma^{2}\kappa}{\mu},\frac{432\overline{\sigma}^{2}\kappa^{2}}{\mu},\frac{12\kappa\lambda\sqrt{10}}{1-\lambda},\frac{5184\overline{\sigma}^{2}\lambda^{2}\kappa^{2}}{\mu(1-\lambda)},\frac{3+\lambda}{1-\lambda}\right\}$, with $\nu=\min\left\{1,\frac{\mu}{432\sigma^{2}\kappa^{2}},\frac{\mu}{72\kappa}\right\}$, we then have, for any $\delta\in(0,1)$ and $T\geq 1$, with probability at least $1-\delta$
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}\kappa\big(1+\log(2d)\big)/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{(2+t_{0})^{3}\Delta_{f}}{n(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
@@
$\times\exp\bigg(\frac{3\alpha_{t}^{2}\nu^{2}n(t+t_{0}+1)^{2}\sigma^{2}\|\nabla f(\overline{x}^{t})\|^{2}}{2}+15\alpha_{t}^{2}\nu\sigma^{2}(t+t_{0}+1)L\big(1+\log(2d)\big)\bigg)$
@@
where in $(a)$ we used the fact that $F_{t}$ and $\|{\mathbf{x}^{t}}-\overline{{\mathbf{x}}}^{t}\|^{2}$ are $\mathcal{F}_{t}$-measurable, $(b)$ follows from Proposition 4, in $(c)$ we use Lemma 2, Proposition 1 and impose the condition $\alpha_{t}\leq\frac{1}{\sigma\sqrt{30(t+t_{0}+1)L}}$, while $(d)$ follows from Proposition 7 and the definition of $F_{t}$, with $b_{t}=\Big(1-\alpha_{t}\mu+3\alpha_{t}^{2}\nu(t+t_{0}+1)L\Big)\frac{t+t_{0}+1}{t+t_{0}}$, $c_{t}=\frac{\alpha_{t}(t+t_{0}+1)L^{2}}{2}$ and $d_{t}=15\alpha_{t}^{2}\sigma^{2}(t+t_{0}+1)L\big(1+\log(2d)\big)$. Taking the full expectation and applying Proposition 4, we get
@@
where the first inequality follows from $\alpha_{t}\leq\frac{1}{\mu}$. Using (43) and (44) in (G), it follows that
@@
where $N_{t}\coloneqq\frac{4a^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+\nicefrac{{9L^{2}}}{{\mu^{2}}})}}{{(1-\lambda)}})}{(1-\lambda)(t+t_{0})^{2}}+\frac{32a^{4}n\sigma^{2}\lambda^{2}L^{2}}{(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{20a^{2}\lambda^{2}L^{2}(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}(t+t_{0})^{8}}$. Define $G_{1}\coloneqq\frac{1080\kappa\sigma^{2}(1+\log(2d))}{\mu}$, $G_{2}\coloneqq\frac{1296\kappa^{2}\lambda^{2}(n\sigma^{2}+\nicefrac{{5\|\nabla\mathbf{f}^{\star}\|^{2}(1+9\kappa^{2})}}{{(1-\lambda)}})}{\mu(1-\lambda)}$, $G_{4}\coloneqq\frac{6480\kappa^{3}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}}$ and $G_{3}\coloneqq\frac{373248n\kappa^{4}\sigma^{2}\lambda^{2}}{\mu(1-\lambda)^{2}}$, and plug into (G), to get
@@
$\mathbb{E}\big[\exp\big(\nu F_{t+1}\big)\big]$ $\leq\exp\bigg(\frac{n(t_{0}+1)^{3}\nu\Delta_{f}}{(t+1+t_{0})^{2}}+4\nu G_{1}+\frac{4\nu G_{2}}{t+1+t_{0}}\bigg)$
@@
Using (G), it can be readily verified that choosing
$\epsilon_{t}^{1}$ $=\frac{\nu^{-1}\log(\nicefrac{{1}}{{\delta}})+4G_{1}}{n(t+t_{0})}+\frac{4G_{2}}{n(t+t_{0})^{2}}+\frac{(t_{0}+1)^{3}\Delta_{f}}{(t+t_{0})^{3}}+\frac{4G_{3}\log(t+t_{0})+4G_{4}/5(t_{0}+1)^{5}}{n(t+t_{0})^{3}},$ (47)
@@
$\frac{1}{n}\sum_{i\in[n]}\big(f(x^{t}_{i})-f^{\star}\big)=\mathcal{O}\bigg(\frac{\nu^{-1}\log(\nicefrac{{2}}{{\delta}})+\sigma^{2}\kappa\big(1+\log(2d)\big)/\mu}{n(t+t_{0})}+\frac{\kappa\lambda^{2}(1+L)(n\sigma^{2}+\|\nabla\mathbf{f}^{\star}\|^{2}\nicefrac{{(1+9\kappa^{2})}}{{(1-\lambda))}}}{\mu(1-\lambda)n(t+t_{0})^{2}}$
$+\frac{(1+t_{0})^{3}\Delta_{f}}{(t+t_{0})^{3}}+\frac{\kappa^{3}\sigma^{2}\lambda^{2}(\kappa\log(t+t_{0})+1)}{\mu(1-\lambda)^{2}(t+t_{0})^{3}}+\frac{\kappa^{3}\lambda^{2}L(1+t_{0})\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{3}}+\frac{\kappa^{2}\lambda^{2}L(1+t_{0})^{6}\|{\mathbf{x}}^{1}-{\mathbf{x}^{\star}}\|^{2}}{(1-\lambda)^{2}n(t+t_{0})^{8}}\bigg).$
@@
Appendix H Numerical Experiments
In this section we perform numerical experiments, with the aim of validating our theory.
@@ Appendix H Numerical Experiments
We consider an instance of (2), with $f_{i}(x)=\frac{1}{2}x^{\top}A_{i}x+b_{i}^{\top}x$, where $A_{i}\in\mathbb{R}^{d\times d}$ is positive definite, making each $f_{i}$ strongly convex, with unbounded gradients. We consider an undirected communication network $G=(V,E)$, corresponding to a randomly generated Erdős–Rényi graph with connectivity parameter $p=0.8$, while the weight matrix $W\in\mathbb{R}^{n\times n}$ is computed using the Metropolis-Hastings weight scheme, e.g., [80]. When queried by user $i\in[n]$ in iteration $t\geq 1$, the $\mathcal{SFO}$ returns a noisy gradient of $f_{i}$ evaluated at $x^{t}_{i}$, i.e., $g^{t}_{i}=A_{i}x^{t}_{i}+b_{i}+z^{t}_{i}$, where $z^{t}_{i}\in\mathbb{R}^{d}$ is a zero-mean Gaussian random vector, making the noise consistent with assumption (A4). We use the DSGD method outlined in Algorithm 1, with the time-varying step-size schedule $\alpha_{t}=\frac{1}{t+1}$ and shared initializatio $x^{1}_{i}=0$, for all $i\in[n]$. We are interested in testing the following two facets of our theory.
@@
Appendix I Deriving the Transient Time